# DMER-Ranker: Learning to Rank Emotion Descriptions in the Absence of Ground Truth

### v2

@@ 4.4 Choice of LLMs
Figure 11: Impact of LLMs. For recognition performance, we report two-class WAF scores.
4.5 Combination of Forward and Reversed Predictions
As demonstrated in Section 4.4, current models demonstrate suboptimal performance regarding multi-run consistency and flip consistency. This observation naturally raises the question: Can we enhance the reliability of the results by aggregating the outputs of multiple rounds of forward and backward inputs? To evaluate this hypothesis, we execute each method twice in forward order and twice in reverse order, generating four preference descriptions $\{p_{1},\cdots,p_{4}\}$. We then employ majority voting to determine the final outcome. Figure 12 presents the impact of this strategy across different MLLMs. Experimental results demonstrate that most models benefit from this approach, including OpenAI GPT and Gemini. However, some exceptions also exist, such as Chat-UniVi and Qwen2.5-VL, indicating that the effectiveness of this strategy is model-dependent.
@@ 4.5 Combination of Forward and Reversed Predictions
4.6 Practical Implementation: DMER-Ranker with Automatic Preference Prediction
This section illustrates our ranking process through a concrete example. We first select $M=10$ models $\mathbf{U}=$Chat-UniVi, LLaMA-VID, mPLUG-Owl, Video-ChatGPT, Otter, SALMONN, Qwen-Audio, Video-LLaVA, VideoChat, VideoChat2} and conduct preference comparisons across $N=332$ samples. This process requires a total of $C(M,2)\times N=14,940$ comparisons. To optimize evaluation efficiency, we apply the strategy described in Section 4.3, utilizing the top four performing models $\mathbf{V}=$AffectGPT, Qwen2.5-VL, Qwen2.5-Omni, Gemini-1.5-Flash} for automated preference prediction. Notably, Gemini-1.5-Flash is the only closed-source model in this group. To minimize evaluation costs, we limit automatic preference prediction to the remaining three open-source models. In this example, we deliberately use different MLLMs for ranking and preference prediction, i.e., $\mathbf{U}\cap\mathbf{V}=\emptyset$.
Figure 13(a) summarizes the final preference prediction results. We then transform these predictions into the matrix $\mathbf{W}$, yielding the results shown in Figure 13(b). Applying the Bradley-Terry algorithm to matrix $\mathbf{W}$, we derive the advantage score $\theta$ for each model. Based on these scores, the models are ranked in descending order as follows: Chat-UniVi, LLaMA-VID, mPLUG-Owl, Video-ChatGPT, Otter, SALMONN, Qwen-Audio, Video-LLaVA, VideoChat, and VideoChat2.
(a) Raw Results
(b) Matrix $\mathbf{W}$
Figure 13: Visualization of raw results and matrix $\mathbf{W}$. a) We present the pairwise comparison results between all models, reporting the number of wins (W), losses (L), and ties (T) for each model pair. b) In this matrix, 1 indicates that model $m_{i}$ wins against model $m_{j}$, 0 indicates that model $m_{i}$ loses to model $m_{j}$, and -1 indicates that no direct comparison is performed between them.
5 Discussion: DMER-Ranker with Manual Preference Annotations
As illustrated in Figure 10, the automatic strategy achieves a maximum accuracy of only approximately 70%, highlighting its inherent limitations for preference prediction. Consequently, human annotation remains indispensable for obtaining reliable ranking results. To balance accuracy requirements with evaluation costs, we propose two practical strategies for human annotation. 1) Reduce the number of comparison samples; 2) Implement a hierarchical comparison strategy. For example, we can divide $M$ models into two subsets, conduct pairwise comparisons within each subset to identify the top-performing model from each, and then perform a final comparison between the two subset winners. This hierarchical approach significantly reduces the total number of required comparisons while maintaining the reliability of the final ranking. Further experiments involving human annotations will be explored in our future work.
6 Conclusion
This paper proposes DMER-Ranker, a novel ranking strategy for the DMER task that eliminates the need for ground-truth descriptions. To achieve this, we shift from the traditional “prediction-ground truth” comparison paradigm into the “prediction-prediction” comparison paradigm, subsequently employing the Bradley-Terry algorithm to derive model rankings from pairwise comparison outcomes. To enhance evaluation efficiency, we explore the potential of automatic preference prediction and introduce DMER-Preference, the first preference dataset focused on human emotions. Extensive experiments across multiple MLLMs demonstrate that current models struggle with accurate preference predictions. Although model-based crowdsourcing improves performance, the predictions still show limited consistency with human annotations. Therefore, at the current stage, automatic evaluation should be treated as a reference rather than a definitive measure. For reliable ranking results, human annotation remains essential.

### v3

@@ Appendix A Impact of External LLMs on S3/S4
Figure 12: Impact of external LLMs. We report two-class WAF scores to evaluate recognition performance. In this table, we reveal the impact of external LLMs on AffectGPT.
Appendix B Combination of Forward and Reversed Predictions
As demonstrated in Section A, current models demonstrate suboptimal performance regarding multi-run consistency and flip consistency. This observation naturally raises the question: Can we enhance the reliability of the results by aggregating the outputs of forward and backward inputs? To evaluate this hypothesis, we execute each method twice in forward order and twice in reverse order, generating four preference descriptions $\{p_{1},\cdots,p_{4}\}$. We then employ majority voting to determine the final outcome. Figure 13 presents the impact of this strategy across different MLLMs. We observe that most models benefit from this approach, including OpenAI GPT and Gemini. However, some exceptions also exist, such as Chat-UniVi and Qwen2.5-VL, indicating that the effectiveness of this strategy is model-dependent.